4,5), AN Jorge Company bottles and distributes B-Lite, a diet soft drink. The beverage is sold for 50 cents per 16-ounce bottle to retailers. For the year 2020, management estimates the following revenues and costs. $1,800,000 Selling expenses-variable $70,000 Sales Direct materials Direct labor Manufacturing overhead- 65,000 20,000 60,000 30,000 Selling expenses-fixed 360,000 Administrative expenses- variable variable 380,000 Administrative expenses- Manufacturing overhead fixed fixed 280,000 Instructions a. Prepare a CVP income statement for 2020 based on managements estimates. (Show column for total amounts only.) b. Compute the break-even point in () units and (2) dollars c. Compute the contribution margin ratio and the margin of safety ratio. (Round to nearest full percent.) d. Determine the sales dollars required to earn net income of $180,000.
Answer #1
Part-1
JORGE COMPANY
CVP Income Statement (Estimated)
Total Per Unit
Sales $18,00,000 $0.50
Less: Variable Expenses :
Direct materials $4,30,000 $0.12
Direct labor $3,60,000 $0.10
Manufacturing overhead—variable $3,80,000 $0.11
Selling expenses—variable $70,000 $0.02
Administrative expenses—variable $20,000 $0.01
Total Variable Costs $12,60,000 $0.35
Contirbution Margin $5,40,000 $0.15
Less: Fixed Expenses :
Manufacturing overhead—fixed $2,80,000
Selling expenses—fixed $65,000
Administrative expenses—fixed $60,000
Total Fixed Expenses $4,05,000
Net Income $1,35,000
*Units = $1800000 /$0.50= 36,00,000
Part-2
Variable cost per bottle= $0.35
Paart-3
Calculation of Break even points:
Total Fixed Expenses (A) $4,05,000
Contirbution Margin Per unit (B) $0.15
Break even point (Units) C =A/B $27,00,000
Sales Price Per unit (D) $0.50
Break even point ($) =C*D $13,50,000
Part-4
Calculation of contribution margin ratio and the margin of safety ratio
Contirbution Margin Per unit (A) $0.15
Sales Price Per unit (B) $0.50
Contribution margin ratio =A/B 30%
Break even point ($) (A) $        1,350,000
Actual Sales (B) $        1,800,000
Margin of safety ratio = (B-A)/B 25%
Part5
Calculation of sales dollars required to earn net income of $180,000.
Total Fixed Expenses (A) $            405,000
Required Profits (B) $            180,000
Contirbution Margin Per unit C) $                   0.15
Required Sales(Units) D =(A+B)/C            3,900,000
Sales Price Per unit (D) $                   0.50
Required Sales ($) =C*D $        1,950,000

Note : Kindly mark positive rating it will help me lot.